Given that the equation normal to the rectangular hyperbola xy = c^2 at the point P(ct, c/t) is t^3*x - ty = c(t^4 - 1)
The normal at P on the hyperbola meets the x - axis at Q and the tangent at P meets the y - axis at R. Show that the locus of the midpoint of QR, as P varies, is
2c^2*xy + y^4 = c^4
1
Expert's answer
2021-02-04T01:36:35-0500

The equation of normal at the point Q:

"t^3x=c(t^4-1)"

Then point Q: "(\frac{c(t^4-1)}{t^3},0)"


The equation of the tangent at the point P:

"y-y_P=f'(P)(x-x_P)"

"y'(x)=-\frac{c^2}{x^2}"

"f'(P)=-\frac{c^2}{c^2t^2}=-\frac{1}{t^2}"

Then:

"y-c/t=-\frac{1}{t^2}(x-ct)"

At the point R:

"y-c\/t=c\/t"

Point R: "(0, 2c\/t)"


Midpoint of QR: "(\frac{c(t^4-1)}{2t^3}, \frac{c}{t})"

Then, for midpoint of QR:

"2c^2xy + y^4=\frac{c^4(t^4-1)}{t^4}+\frac{c^4}{t^4}=\frac{c^4t^4}{t^4}=c^4"